Circular No. 7109

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ab IN ANONYMOUS GALAXY M. Schwartz, Cottage Grove, OR, reports his discovery of a supernova

(mag 18.3) on a CCD image taken on Feb. 8.5074 UT with the Tenagra I internet-controlled 0.35-m supernova patrol telescope in Sonoita, AZ. SN 1999ab is located at α = 10^{h}25^{m}00^{s}.64, δ = +53^{o}52′37′′.0 (equinox 2000.0), which is 4′′.8 east and 6′′.6 north of the host galaxy’s center.

A. Goobar, Stockholm University; I. M. Hook, Royal Observatory, Edinburgh; and G. Aldering and P. Nugent, Lawrence Berkeley National Laboratory, report that CCD spectra taken with the Nordic Optical Telescope on Feb. 13 UT show that SN 1999ab (at V = 18.5) is a type II-supernova ∼ 1–2 months after maximum light. Strong Hβ and Hα lines are present. The redshift of the host galaxy was determined through Hα, Hβ, and O III lines, placing it and the supernova at 9400 km/s. This supernova was also confirmed with photometric U,B,V,R images taken by E. Gates at the Lick 1.0-m telescope on Feb. 13.

SUPERNOVA 1999aa IN NGC 2595 Q. Y. Qiao, J. Y. Wei, Y. L. Qiu, and J. Y. Hu, Beijing Astronomical

Observatory (BAO), report their independent discovery of SN 1999aa (cf. IAUC 7108) during the course of the BAO supernova survey with the 0.6-m telescope at Xinglong station, providing a position of α = 8^{h}27^{m}42^{s}.11, δ = +21^{o}29′15′′.5 (equinox 2000.0), which is ∼ 1′′.4 east and 31′′.0 north of the nucleus of NGC 2595. Unfiltered CCD images taken on Feb. 11.6 and 12.6 UT show SN 1999aa at mag ∼ 16.3 and 16.0, respectively, but it was not present on similar images obtained on Jan. 15 and Feb. 5.

S. Nakano, Sumoto, Japan, reports the independent discovery of SN 1999aa by R. Kushida (Yatsugatake South Base Observatory) at V = 15.5 on Feb. 13.663 UT on a CCD image (limiting mag 19) obtained with a 0.40m f/6.3 Schmidt-Cassegrain reflector, from which Y. Kushida measured the following position end figures for the supernova: 42^{s}.13, 15′′.5 (or 1′′.7 east and 30′′.8 north of the galaxy’s center). R. Kushida notes that dozens of past exposures taken between 1995 and 1999 show nothing at this position, the most recent being obtained on 1999 Jan. 21 (limiting mag 17.5).

M. Villi, Forlí, Italy, reports that an unfiltered CCD image (limiting mag 18) obtained on Feb. 6.91 UT shows nothing at the position of SN 1999aa.

1999 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green